Assume the population is normally distributed with: 

Xbar = 96.44           S =10.2        n=15                 t=1.76

Construct a​ 90% confidence interval estimate for the population​ mean, μ


1
Expert's answer
2021-04-15T07:40:52-0400

"c = 0.90 \\

n = 15 \\

\bar{x} = 96.44 \\

t_{α/2}= 1.76"

The margin of error is:

"E = t_{α/2} \times \frac{S}{\sqrt{n}} = 1.76 \times \frac{10.2}{\sqrt{15}} = 4.63"

The confidence interval:

"\bar{x} -E < μ < \bar{x} +E \\

96.44 -4.63 < μ < 96.44 +4.63 \\

91.81 < μ < 101.07"